Sumter Corporation uses the weighted-average method in its process costing system.

The following data pertain to operations in the first processing department for a recent month: Work in process, beginning: Units in process 6,000 Percent complete with respect to materials 60 % Percent complete with respect to conversion 20 % Costs in the beginning inventory: Materials cost $ 78,200 Conversion cost $ 3,600 Units started during the month ? Units completed and transferred out during the month 70,000 Costs added to production during the month: Materials cost $ 286,600 Conversion cost $ 216,000 Work in process, ending: Units in process 8,000 Percent complete with respect to materials 75 % Percent complete with respect to conversion 25 % a. Determine the equivalent units of production. b. Determine the costs per equivalent unit. c. Determine the cost of ending work in process inventory. d. Determine the cost of the units transferred to the next department.
Answer #1

a. Equivalent units of production.

Materials: 76,000 ; Conversion: 72,000

Working:

Beginning WIP + Units started = 6,000+72,000 = 78,000

Materials

Conversion

Units out

70,000

70,000

70,000

Ending WIP (75%, 25%)

8,000

6,000

2,000

78,000

76,000

72,000

?

b. Cost per equivalent unit

Materials: $4.8 ; Conversion: $3.05

Working:

Materials

Conversion

Beginning WIP

78,200

3,600

Cost added during the year

286,600

216,000

Total amount

364,800

219,600

?

Materials: 364,800 / 76,000 = 4.8

Conversion: 219,600 / 72,000 = 3.05

?

c. Cost of ending work in process inventory: $34,900

Working: 6000 X $4.8 + 2000 X $3.05 = $34,900

?

d. Cost of the units transferred to the next department: $549,500

Working: 70,000 X $4.8 + 70,000 X $3.05 = $549,500

?

?